OpenStudy (anonymous):

The concentration C of a certain chemical in the bloodstream t hours after injection into muscle tissue is given by C=(3t)/(27+t^3) When is the concentration greatest?

OpenStudy (anonymous):

Answer: 2.381 hours

OpenStudy (anonymous):

@yun2thejae

OpenStudy (anonymous):

Oh I guess it is asking you to find the absolute maximum of that graph. give me second. I'm a tad rusty .

OpenStudy (anonymous):

Here are the steps: 1. Take the derivative of that equation. 2. Set the derivative equal to zero to find critical numbers. 3. Test the critical numbers to find local maximum points. 4. Plug those maximum points back into the original equation. 5. The largest number is your answer. in this case, 2.381.

OpenStudy (anonymous):

I'm going to try to work this out,. but you try it also in case I do it too slow.

OpenStudy (anonymous):

u did it already or was that a guess

OpenStudy (anonymous):

? you gave me the answer up there. I have not finished it yet, I'm haing a tad of trouble finding the critical numbers to test

OpenStudy (anonymous):

Oh wait, I see why I am having problems. give me a second.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Take the derivative, you get (81+3t^3-9t^3)/(27+t^3)^2. set that equal to zero. you have to solve for 81+3t^3-9t^3=0

OpenStudy (anonymous):

81=6t^3 13.5=t^3 Cube root of 13.5 = t.

OpenStudy (anonymous):

if you test that point, it should be the absolute maximum of the equation. cube root of 13.5 equals the answer you gave me up there.

OpenStudy (anonymous):

sorry it took so long. I have the lights turned off while trying to hand write math LOL. I kept messing up simple algebra.

OpenStudy (anonymous):

i got \[\frac{ 81-6t^3 }{ (27+t^3)^2 }\] for my derivative

OpenStudy (anonymous):

is that 81-6t^3? I can't see very well

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

yeah that is correct. now set it equal to zero. When you do that, the bottom goes away because you cross multiplied it with the zero. so you end up with 81-6t^3=0

OpenStudy (anonymous):

oh okay ty o.O

OpenStudy (anonymous):

solve for t and that is your answer. Ideally, you're expected to check to see if it actually a maximum, but in this case there is only one number so..you don't have to. Although, I would on the Ap exam though.

OpenStudy (anonymous):

thx :P

OpenStudy (anonymous):

You're welcome. Anything else?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Post it up and i'll see if I can help.